Mathematics 243 – Algebraic Structures Definitions Quiz 4 – October 5, 2012

Your Name:

A) (5) What does it mean for f : A → B to be a one-to-one mapping?

B) (5) Let R be the relation on Z defined by xRy ⇔ x = ky for some integer k. Is R an equivalence relation? Why or why not?